# Unleashing the Power of Task-Specific Directions in Parameter Efficient Fine-tuning

### v1

@@ C.2 Why should TSD be Established on $\mathbf{W}$?
C.3 Why Must TSD be in the Core Bases Rather than Other Global Bases?
@@
Firstly, if TSD were defined based on the change rates, since the coordinate values of $\mathbf{W}$ on the other global bases are zero, even a minimal projection of $\Delta\mathbf{W}$ onto these global bases could result in an excessively high change rate, which would be unreasonable. Secondly, if TSD were defined based on the absolute magnitude of coordinate changes, consider a scenario where a coordinate value in the core basis and a coordinate value in the global basis change by the same amount. If the original coordinate value in the core basis was significantly larger than the change, the change, even if substantial, could be viewed as a minor perturbation relative to the original value. Therefore, the same magnitude of change does not necessarily imply that the corresponding directions are equally important. For these reasons, we define TSD based solely on the rate of change and restrict their definition to the core bases.
Appendix D Details on LoRA-Dash
D.1 Algorithm of LoRA-Dash
@@
D.2 Matrix Operation of Dash Phase
@@
Appendix E Details on Experiments
E.1 Implementation Details
We compare LoRA-Dash mainly with LoRA. The ranks of LoRA-Dash, LoRA and other methods are varied among {2, 4, 8, 16, 32, 64}. All methods are implemented using the publicly available PyTorch Paszke et al. (2019) framework, and all experiments are conducted on NVIDIA A100 GPUs [^5] We have also test that most of the experiments can also be conducted on one consumer GPU resource such as NVIDIA RTX3090.. The hyper-parameter $t$ of LoRA-Dash is set to 100, and $s=8$. We fine-tune all the layers of different models in all the experiments, and the detailed experimental settings are shown in the following Secs. E.4-E.6.
E.2 Details on Fully Fine-tuning LLaMA in Section 3.3
To further explore the properties of TSD, we fully fine-tune LLaMA-7B Touvron et al. (2023a) on commonsense reasoning tasks. For the commonsense reasoning tasks, which consist of 8 distinct sub-tasks, each defined with specific training and testing sets, we have adopted the approach in Hu et al. (2023). Following this method, we amalgamated the training datasets from all sub-tasks to construct a comprehensive final training dataset. More details can be found in Sec. E.4. The experiment was conducted on 8 A100 GPUs.
@@ E.1 Implementation Details
E.3 Details on Fine-tuning LLaMA with LoRA in Section 3.4
Figure 7: We track the recall of TSD every 100 training steps in the query, key and value layers of the LLaMA-7B during LoRA fine-tuning, analyzing how well the continuous updated $\Delta\mathbf{W}$ captures TSD. The left and right figure settings are the same as those in Fig. 2.
We fine-tune LLaMA-7B using LoRA under three ranks, and for every 100 training steps we record the top 8 core directions of $\mathbf{W}$ which exhibit the highest change rates based on the projection of continuously updated $\Delta\mathbf{W}$ (i.e., the directions related to top 8 largest $\delta_{i}=\mathbf{u}_{i}^{\mathsf{T}}\Delta\mathbf{W}\mathbf{v}_{i}/\sigma_{i}$, where $\mathbf{u}_{i}$, $\mathbf{v}_{i}$ and $\sigma_{i}$ are the corresponding left/right singular vectors and values of $\mathbf{W}$). These directions serve as the “TSD” predicted by LoRA. We denote these directions as “launched TSD” (LTSD) for differentiation. We then identify the top 4 or 16 TSD with the highest change rates in Sec. 3.3.
@@
    How many of the top 4 TSD are contained in the 8 LTSD;
@@
    How many LTSD are contained in the top 16 TSD.
The purpose of the first experiment is more intuitive, aiming to confirm whether the TSD with the top highest change rates can indeed be captured by $\Delta\mathbf{W}$ from LoRA. The second experiment, which may seem a bit more obscure, is designed to address concerns that, besides capturing the top-ranked TSD, the LTSD identified might include irrelevant directions. Therefore, this second setup allows us to examine whether the LTSD predominantly fall within the upper echelons of TSD based on their change rates, providing insight into the overall relevance and accuracy of the LTSD identified by LoRA. Specifically, we employ two metrics to quantify the quality of LTSD:
  • Accuracy: Measures how many of the 8 LTSD are correctly included in the top 16 TSD.
  • Recall: Assesses how many of the top 4 TSD are captured among the 8 LTSD.
The results presented in Figs 2 and 7 lead to two significant observations regarding the effectiveness of using $\Delta\mathbf{W}$ to identify TSD:
  • Precision in Identifying Crucial TSD: The LTSD captured by $\Delta\mathbf{W}$ include several of the actual TSD with the highest change rates, demonstrating that the predicted directions are not only accurate but pinpoint some of the most critical TSD.
  • Overall Coverage of Significant TSD: The entirety of the LTSD resides within the upper echelon of TSD in terms of change rates. This suggests that while the primary directions identified are indeed the most vital, the remaining directions predicted also belong to the upper tier of important TSD.
These insights affirm the utility of $\Delta\mathbf{W}$ in accurately capturing and representing the information of real TSD. The method shows potential in leveraging learned directional changes to identify and utilize key features specific to given tasks, underscoring the robustness of $\Delta\mathbf{W}$ as a predictive tool in practical applications.
For implementation details, please refer to Sec. E.4.
@@
E.4 Details on Commonsense Reasoning Task
@@
E.5 Details on Natural Language Understanding Task
@@
E.6 Details on Subject-driven Generation Task
@@
E.7 Details on the Amplification of TSD in Sec. 6.2
To explore to what extent LoRA-Dash can amplify the information of the TSD it determines, we project the pretrained weights $\mathbf{W}$ and the weights merged after training, i.e., $\mathbf{W}+\Delta\mathbf{W}_{all}$, onto the LTSD. Let $\bar{\mathbf{U}}=\begin{bmatrix}\bar{\mathbf{u}}_{1}&\ldots&\bar{\mathbf{u}}_{s}\end{bmatrix}$ and $\bar{\mathbf{V}}=\begin{bmatrix}\bar{\mathbf{v}}_{1}&\ldots&\bar{\mathbf{v}}_{s}\end{bmatrix}$, we can then define the amplification factor of $\Delta\mathbf{W}_{all}$ as the Frobenius norms of these two weights on these directions, which is given by:
@@
Moreover, as shown in Sec. C.2, TSD are indeed the projection of $\Delta\mathbf{W}^{*}$’s direction onto the core bases of $\mathbf{W}$. Therefore, we also explored the amplification factor along the direction of $\Delta\mathbf{W}_{all}$. We replaced $\bar{\mathbf{U}}$ and $\bar{\mathbf{V}}$ in the previous amplification factor calculation with the top $r$ left and right singular vectors of $\Delta\mathbf{W}_{all}$, where $r$ is the rank setting of LoRA-Dash. We reported the corresponding results in Table. 18, where we can draw a conclusion:
Table 18: Amplification of $\Delta\mathbf{W}_{all}$ of LoRA-Dash onto TSD or $\Delta\mathbf{W}_{all}$’s direction. We report the amplification factor from the self-attention key (k) projection of the first, middle and last layer of LLaMA-7B, as well as the average of all q, k, and v modules.
@@
  • The amplification factor on the directions of $\Delta\mathbf{W}_{all}$ is significantly higher than that on the TSD. Since TSD can be considered as a projection of $\Delta\mathbf{W}$ onto the coordinate system of $\mathbf{W}$, it is natural that some information is lost in this projection for we only select $s$ TSD, leading to a lower amplification factor on the TSD compared to before the projection.
Additionally, we observed an interesting trend: the amplification factor on $\Delta\mathbf{W}_{all}$’s directions significantly decreases as the rank increases. In LoRA Hu et al. (2021), the authors claimed that the occurrence of a small number of TSD leads to this situation. However, a more plausible explanation is that higher rank corresponds to $\Delta\mathbf{W}$ containing more irrelevant directions, thereby reducing the amplification factor.
To validate this, consider the setting of $r=32$ of $\Delta\mathbf{W}_{all}$. The 32 directions of $\Delta\mathbf{W}_{all}$ include some irrelevant directions, which is why the amplification factor is lower. This implies that if the irrelevant directions among these 32 are removed, the amplification factor could potentially increase. To explore whether this is the case, we recalculated the amplification factors using the top 2, 4, 8 and 16 singular directions of $\Delta\mathbf{W}_{all}$, and the results are shown in Table. 19. The observation that the amplification factor decreases as projections are made onto an increasing number of the top singular directions of $\Delta\mathbf{W}_{all}$ indicates that $\Delta\mathbf{W}_{all}$’s singular directions do indeed contain irrelevant directions in its bottom singular directions. It is not surprising, since TSD are only a few directions, as shown in its properties. Therefore, it also motivates us to select several top directions in pre-launch phase of LoRA-Dash.
@@
Appendix F More Experiment Results
F.1 More Results for Subject-driven Generation Task
@@
F.2 More Results for the Alignment on the Directions of LoRA-Dash
@@
Table 20: Alignment on the directions of LoRA-Dash with TSD. We report the percentage from the self-attention value (v) projection of the first, middle and last layer of two models, as well as the average percentage of all q, k, and v modules.
@@
Table 21: Alignment on the directions of LoRA-Dash with TSD. We report the percentage from the self-attention value (v) projection of the first, middle and last layer of two models, as well as the average percentage of all q, k, and v modules.
@@
F.3 More Results for Ablation Study
@@
Indeed, fine-tuning leverages a model’s broad pre-trained knowledge base, adapting it to learn specific, nuanced details relevant to particular downstream tasks. The essence of TSD lies in their representation of the singular directions where the original model weights exhibit the most significant changes during adaptation to new tasks. Intuitively, these TSD, being the most altered, should indeed offer the most impactful directions for adjustment during fine-tuning, potentially yielding the best performance.
Analyzing the properties of TSD reveals that they typically reside along the directions corresponding to smaller, yet not the smallest, singular values. If we choose to dash in the directions associated with the smallest singular values—the “bottom” directions—it’s plausible that some of these might overlap with the TSD. Hence, dashing in the bottom directions might produce results inferior to directly using the TSD.
@@
Lastly, a random selection of directions introduces unpredictability in outcomes. However, there is a chance that such random choices might inadvertently select appropriate directions or even TSD, thus performing better than top selections but generally worse than bottom choices due to the randomness. This method could strike a balance between leveraging underlying model strengths and adapting to new specifics, albeit inconsistently.
In our empirical tests, the results largely align with our predictions. However, there are some exceptions where the effectiveness of top directions surpasses random or bottom, or where the bottom directions approach or even slightly exceed the performance of the TSD. Please look for a counterexample in the Fig. 9(a). We consider these variations to be within the expected range, acknowledging that it’s unrealistic to expect uniform trends across all tasks and models. The trends we have identified represent a more general, broadly applicable pattern.
@@
F.4 Are TSD “Task-specific”?
We sought to verify whether TSD are genuinely task-specific, that is, whether the most critical directions differ for each task. To investigate this, we employed a prototypical example: the subject-driven generation task. In this setting, we tracked the TSD captured by LoRA-Dash when different subjects served as targets and analyzed the relationships between these TSD. The results are presented in the Fig. 10 (Please refer to the caption of Fig. 10 for detailed information).
Figure 10: The results of the TSD similarity between different tasks (subjects). For the first row: Images from different subjects. The second and third rows contain results extracted from two different layers of U-net Ronneberger et al. (2015). Each subplot illustrates the similarity of TSD captured for two different subjects. The title of each subplot indicates the respective subjects, with the directional order of the first subject represented by columns and the second by rows. The deep blue blocks highlight where the directions for the two subjects coincide.
The experimental results affirm the task-specific nature of TSD; indeed, each task exhibits distinct TSD with minimal overlap between tasks. Moreover, even when tasks share the same direction, its importance varies significantly between them. For instance, as depicted in the central subfigure of the first column in Fig. 10, both “Dog A” and “Cat” share a common direction, yet it ranks as the seventh most significant for “Dog A” and the second most significant for “Cat”. Furthermore, tasks that are closely related tend to share more directions, as observed between “Dog A” and “Dog B”; however, the importance of these shared directions still differs for each task. This variability underscores the fundamentally task-specific property of TSD, illustrating their unique and variable impact across different contexts.
F.5 Convergence Curve

### v2

@@ B.2 Why should TSD be Established on $\mathbf{W}$?
B.3 Why Must TSD be in the Core Bases Rather than Other Global Bases?
@@
Firstly, if TSDs were defined based on the change rates, since the coordinate values of $\mathbf{W}$ on the other global bases are zero, even a minimal projection of $\Delta\mathbf{W}$ onto these global bases could result in an excessively high change rate, which would be unreasonable. Secondly, if TSDs were defined based on the absolute magnitude of coordinate changes, consider a scenario where a coordinate value in the core basis and a coordinate value in the global basis change by the same amount. If the original coordinate value in the core basis was significantly larger than the change, the change, even if substantial, could be viewed as a minor perturbation relative to the original value. Therefore, the same magnitude of change does not necessarily imply that the corresponding directions are equally important. For these reasons, we define TSDs based solely on the rate of change and restrict their definition to the core bases.
Appendix C Details on LoRA-Dash
C.1 Algorithm of LoRA-Dash
@@
C.2 Matrix Operation of Dash Phase
@@
Appendix D Details on Experiments
D.1 Implementation Details
We compare LoRA-Dash mainly with LoRA. The ranks of LoRA-Dash, LoRA and other methods are varied among {2, 4, 8, 16, 32, 64}. All methods are implemented using the publicly available PyTorch Paszke et al. (2019) framework, and all experiments are conducted on NVIDIA A100 GPUs [^4] We have also test that most of the experiments can also be conducted on one consumer GPU resource such as NVIDIA RTX3090.. The hyper-parameter $t$ of LoRA-Dash is set to 100, and $s=8$. We fine-tune all the layers of different models in all the experiments, and the detailed experimental settings are shown in the following Secs. D.4-D.6.
D.2 Details on Fully Fine-tuning LLaMA in Section 3.3
To further explore the properties of TSDs, we fully fine-tune LLaMA-7B Touvron et al. (2023a) on commonsense reasoning tasks. For the commonsense reasoning tasks, which consist of 8 distinct sub-tasks, each defined with specific training and testing sets, we have adopted the approach in Hu et al. (2023). Following this method, we amalgamated the training datasets from all sub-tasks to construct a comprehensive final training dataset. More details can be found in Sec. D.4. The experiment was conducted on 8 A100 GPUs.
@@ D.1 Implementation Details
D.3 Details on Fine-tuning LLaMA with LoRA in Section 3.4
Figure 7: We track the recall of TSDs every 100 training steps in the query, key and value layers of the LLaMA-7B during LoRA fine-tuning, analyzing how well the continuous updated $\Delta\mathbf{W}$ captures TSDs. The left and right figure settings are the same as those in Fig. 2.
We fine-tune LLaMA-7B using LoRA under three ranks, and for every 100 training steps we record the top 8 core directions of $\mathbf{W}$ which exhibit the highest change rates based on the projection of continuously updated $\Delta\mathbf{W}$ (i.e., the directions related to top 8 largest $\delta_{i}=\mathbf{u}_{i}^{\mathsf{T}}\Delta\mathbf{W}\mathbf{v}_{i}/\sigma_{i}$, where $\mathbf{u}_{i}$, $\mathbf{v}_{i}$ and $\sigma_{i}$ are the corresponding left/right singular vectors and values of $\mathbf{W}$). These directions serve as the “TSDs” predicted by LoRA. We denote these directions as “launched TSDs” (LTSDs) for differentiation. We then identify the top 4 or 16 TSDs with the highest change rates in Sec. 3.3.
@@
    How many of the top 4 TSDs are contained in the 8 LTSDs;
@@
    How many LTSDs are contained in the top 16 TSDs.
The purpose of the first experiment is more intuitive, aiming to confirm whether the TSDs with the top highest change rates can indeed be captured by $\Delta\mathbf{W}$ from LoRA. The second experiment, which may seem a bit more obscure, is designed to address concerns that, besides capturing the top-ranked TSDs, the LTSDs identified might include irrelevant directions. Therefore, this second setup allows us to examine whether the LTSDs predominantly fall within the upper echelons of TSDs based on their change rates, providing insight into the overall relevance and accuracy of the LTSDs identified by LoRA. Specifically, we employ two metrics to quantify the quality of LTSDs:
  • Accuracy: Measures how many of the 8 LTSDs are correctly included in the top 16 TSDs.
  • Recall: Assesses how many of the top 4 TSDs are captured among the 8 LTSDs.
The results presented in Figs 2 and 7 lead to two significant observations regarding the effectiveness of using $\Delta\mathbf{W}$ to identify TSDs:
  • Precision in Identifying Crucial TSDs: The LTSDs captured by $\Delta\mathbf{W}$ include several of the actual TSDs with the highest change rates, demonstrating that the predicted directions are not only accurate but pinpoint some of the most critical TSDs.
  • Overall Coverage of Significant TSDs: The entirety of the LTSDs resides within the upper echelon of TSDs in terms of change rates. This suggests that while the primary directions identified are indeed the most vital, the remaining directions predicted also belong to the upper tier of important TSDs.
These insights affirm the utility of $\Delta\mathbf{W}$ in accurately capturing and representing the information of real TSDs. The method shows potential in leveraging learned directional changes to identify and utilize key features specific to given tasks, underscoring the robustness of $\Delta\mathbf{W}$ as a predictive tool in practical applications.
For implementation details, please refer to Sec. D.4.
@@
D.4 Details on Commonsense Reasoning Task
@@
D.5 Details on Natural Language Understanding Task
@@
D.6 Details on Subject-driven Generation Task
@@
D.7 Details on the Amplification of TSD in Sec. 6.2
To explore to what extent LoRA-Dash can amplify the information of the TSDs it determines, we project the pretrained weights $\mathbf{W}$ and the weights merged after training, i.e., $\mathbf{W}+\Delta\mathbf{W}_{all}$, onto the LTSDs. Let $\bar{\mathbf{U}}=\begin{bmatrix}\bar{\mathbf{u}}_{1}&\ldots&\bar{\mathbf{u}}_{s}\end{bmatrix}$ and $\bar{\mathbf{V}}=\begin{bmatrix}\bar{\mathbf{v}}_{1}&\ldots&\bar{\mathbf{v}}_{s}\end{bmatrix}$, we can then define the amplification factor of $\Delta\mathbf{W}_{all}$ as the Frobenius norms of these two weights on these directions, which is given by:
@@
Moreover, as shown in Sec. B.2, TSDs are indeed the projection of $\Delta\mathbf{W}^{*}$’s direction onto the core bases of $\mathbf{W}$. Therefore, we also explored the amplification factor along the direction of $\Delta\mathbf{W}_{all}$. We replaced $\bar{\mathbf{U}}$ and $\bar{\mathbf{V}}$ in the previous amplification factor calculation with the top $r$ left and right singular vectors of $\Delta\mathbf{W}_{all}$, where $r$ is the rank setting of LoRA-Dash. We reported the corresponding results in Table. 18, where we can draw a conclusion:
Table 18: Amplification of $\Delta\mathbf{W}_{all}$ of LoRA-Dash onto TSDs or $\Delta\mathbf{W}_{all}$’s direction. We report the amplification factor from the self-attention key (k) projection of the first, middle and last layer of LLaMA-7B, as well as the average of all q, k, and v modules.
@@
  • The amplification factor on the directions of $\Delta\mathbf{W}_{all}$ is significantly higher than that on the TSDs. Since TSDs can be considered as a projection of $\Delta\mathbf{W}$ onto the coordinate system of $\mathbf{W}$, it is natural that some information is lost in this projection for we only select $s$ TSDs, leading to a lower amplification factor on the TSD compared to before the projection.
Additionally, we observed an interesting trend: the amplification factor on $\Delta\mathbf{W}_{all}$’s directions significantly decreases as the rank increases. In LoRA Hu et al. (2021), the authors claimed that the occurrence of a small number of TSDs leads to this situation. However, a more plausible explanation is that higher rank corresponds to $\Delta\mathbf{W}$ containing more irrelevant directions, thereby reducing the amplification factor.
To validate this, consider the setting of $r=32$ of $\Delta\mathbf{W}_{all}$. The 32 directions of $\Delta\mathbf{W}_{all}$ include some irrelevant directions, which is why the amplification factor is lower. This implies that if the irrelevant directions among these 32 are removed, the amplification factor could potentially increase. To explore whether this is the case, we recalculated the amplification factors using the top 2, 4, 8 and 16 singular directions of $\Delta\mathbf{W}_{all}$, and the results are shown in Table. 19. The observation that the amplification factor decreases as projections are made onto an increasing number of the top singular directions of $\Delta\mathbf{W}_{all}$ indicates that $\Delta\mathbf{W}_{all}$’s singular directions do indeed contain irrelevant directions in its bottom singular directions. It is not surprising, since TSDs are only a few directions, as shown in its properties. Therefore, it also motivates us to select several top directions in pre-launch phase of LoRA-Dash.
@@
Appendix E More Experiment Results
E.1 More Results for Subject-driven Generation Task
@@
E.2 More Results for the Alignment on the Directions of LoRA-Dash
@@
Table 20: Alignment on the directions of LoRA-Dash with TSDs. We report the percentage from the self-attention value (v) projection of the first, middle and last layer of two models, as well as the average percentage of all q, k, and v modules.
@@
Table 21: Alignment on the directions of LoRA-Dash with TSDs. We report the percentage from the self-attention value (v) projection of the first, middle and last layer of two models, as well as the average percentage of all q, k, and v modules.
@@
E.3 More Results for Ablation Study
@@
Indeed, fine-tuning leverages a model’s broad pre-trained knowledge base, adapting it to learn specific, nuanced details relevant to particular downstream tasks. The essence of TSDs lies in their representation of the singular directions where the original model weights exhibit the most significant changes during adaptation to new tasks. Intuitively, these TSDs, being the most altered, should indeed offer the most impactful directions for adjustment during fine-tuning, potentially yielding the best performance.
Analyzing the properties of TSDs reveals that they typically reside along the directions corresponding to smaller, yet not the smallest, singular values. If we choose to dash in the directions associated with the smallest singular values—the “bottom” directions—it’s plausible that some of these might overlap with the TSDs. Hence, dashing in the bottom directions might produce results inferior to directly using the TSDs.
@@
Lastly, a random selection of directions introduces unpredictability in outcomes. However, there is a chance that such random choices might inadvertently select appropriate directions or even TSDs, thus performing better than top selections but generally worse than bottom choices due to the randomness. This method could strike a balance between leveraging underlying model strengths and adapting to new specifics, albeit inconsistently.
In our empirical tests, the results largely align with our predictions. However, there are some exceptions where the effectiveness of top directions surpasses random or bottom, or where the bottom directions approach or even slightly exceed the performance of the TSDs. Please look for a counterexample in the Fig. 9(a). We consider these variations to be within the expected range, acknowledging that it’s unrealistic to expect uniform trends across all tasks and models. The trends we have identified represent a more general, broadly applicable pattern.
@@
E.4 Are TSDs “Task-specific”?
We sought to verify whether TSDs are genuinely task-specific, that is, whether the most critical directions differ for each task. To investigate this, we employed a prototypical example: the subject-driven generation task. In this setting, we tracked the TSDs captured by LoRA-Dash when different subjects served as targets and analyzed the relationships between these TSDs. The results are presented in the Fig. 10 (Please refer to the caption of Fig. 10 for detailed information).
Figure 10: The results of the TSDs similarity between different tasks (subjects). For the first row: Images from different subjects. The second and third rows contain results extracted from two different layers of U-net Ronneberger et al. (2015). Each subplot illustrates the similarity of TSDs captured for two different subjects. The title of each subplot indicates the respective subjects, with the directional order of the first subject represented by columns and the second by rows. The deep blue blocks highlight where the directions for the two subjects coincide.
The experimental results affirm the task-specific nature of TSDs; indeed, each task exhibits distinct TSDs with minimal overlap between tasks. Moreover, even when tasks share the same direction, its importance varies significantly between them. For instance, as depicted in the central subfigure of the first column in Fig. 10, both “Dog A” and “Cat” share a common direction, yet it ranks as the seventh most significant for “Dog A” and the second most significant for “Cat”. Furthermore, tasks that are closely related tend to share more directions, as observed between “Dog A” and “Dog B”; however, the importance of these shared directions still differs for each task. This variability underscores the fundamentally task-specific property of TSDs, illustrating their unique and variable impact across different contexts.
E.5 Convergence Curve